OpenStudy (anonymous):

who can help me to Solve this problem. Brandon can paint a fence in 12 hours and Elaine can paint the same fence in 11 hours. How long will they take to paint the fence if they work together? Please choose the rihgt answer. a. 5 \frac{13}{24} hr b. 5 \frac{1}{4} hr c. 5 \frac{17}{23} hr d. 11 \frac{1}{2} hr

OpenStudy (anonymous):

1/12+1/11=1/x 11x+12x=132 23x=132 x=5 17/23

OpenStudy (anonymous):

\[\frac{12\times 11}{12+11}\]

OpenStudy (anonymous):

But what is the answer? A, b, c,or d

OpenStudy (anonymous):

x=5 17/23